2012, the imaginary nation of Dorados had a population of 8,000 and real GDP of 3,000,000. During year its real GDP grew by with this growth in real GDP? a. 2% population growth and 6% real GDP growth b, 690 population growth and 2% real GDP growth C. 4% population growth and 7% real GDP growth d, 7% population growth and 4% real GDP growth about 2.9%. Whichofthe following sets of growth rates is consistent 3.300000 op
Answer #1

Real GDP = 3000000

RGDP growth rate = 2.9%

New RGDP = 3000000(1+2.9%) = 3087000

Population =8000

RGDP per capita = RGDP/ Population = 3000000/8000 = 375$ per person

New RGDP per capita= 3087000/8000 = 386$ per person

%Change in RGDP percapita = ((386/300)-1)*100 = 28.67%

Sno % growth in RGDP % Growth in population New RGDP New population RGDP per capita
A 6 2 3000000(1+6%) = 3180000 8160 390
B 2 6 3060000 8480 361
C 7 4 3210000 8320 386
D 4 7 3120000 8560 364

Answer: C